The United States has a population of 327 million and has 206 million people that are of working age within the population. Of those 206 million people, 153.5 million are employed and 6.5 million are unemployed. The equilibrium real wage is $22 (with economy wide price level P = 1). Q 3. Calculate the employment rate and label it on the graph.
Answer #1

Total employed people(E)=153.5million

total population(P)=327million

hence employment rate=E/P=153.5/327=0.4694=46.94%

Give the answer a thumbs up if it helped you.

comment if you have anything to ask regarding the answer.